The Mass spectra of several isomeric bicyclic bromides, C7H9Br are shown. Match each specteum to one of the compounds. Explain.
The mass spectra of several isomeric bicyclic bromides, CHlBr, are shown. Match each spectrum to one of the compounds.Explain your reasoning. Br 100 ??? ?, 10 mle 100 150 mle 80 E 60 240 T 172 mle 100 150 150 50 50 100 T 20 100 150
Answer #1

I = B, II = E, III = D, IV = A, V = C

Among the compounds I, II, III, IV and V.

Only compound “III” don’t have any double bond, means it should give you only even mass on fragmentation, and most intense peak may be C7H9+, m/z = 93, among the spectra it corresponds to “D

Compounds “I” and “II” should give similar fragmentation, most intense peak may be C5H6+, m/z = 66, corresponds to the spectra “B” and “E”, it will be impossible to different the isomer just with mass spectra.

Compounds “IV” and “V” undergoes, similar fragmentation, most intense peak may be C5H5Br+, m/z = 144. But compound “IV” can lose the Br atom, and form C7H9+, m/z = 93. This infer that the spectra of “IV” is “A” and “V” is “C